/3 peiets The paint used to make ines on roads must reflect enough-ght to be clearly visitie at night. Let μ denote the tue average-eflectormeter reading for new type of paint under cons deration A test of Mg: μι-20 versus H > 20 w be based on a random sampie of size from a normal population distnation. Whet concusion is appropniate in each of the follewing situations? (Round your P-values to three decimal places (a) 2 3.3,a-0.0 value State the conclusion in the problem context Do not reject the null hypothess.There is sultficent evidence to conclude that the new paint has a reflecto meter reading higher than 20 Reject the null hypothesis. There is not sufficient evidence to condude that the new paint has a refectometer reading higher than 20 Do not reject the ull hypothess. There is not sulficient evidence to conclude that the new paint has areflectometer reading higher than 20. Reject the null hypothesis. There is sufficient evidence to conclude that the new paint has a reflectometer reading higher than 20 (o) z-1.6,a-0.0 р.vaue- Reject the null hypothesis. There is not sufficiert evidence to condude that the new paint has a reflectometer reading higher than 20. Do not reject the mull hypothess. There is sufficent evidence to conclude that the new paint hes a reflecto meter reading higher than 20. Do not reject the null hypothess.There is not sutficient evidence to conclude that the new paint has a reflectometer reading higher than 20. Reject the null hypothesis. There is sufficient evidence to conclude that the new paint has a reflectometer reading higher than 20 (c) Ζ.-0.5.α-0.10 value State the conclusion in the problem context Do not reject the null hypothess.There is not sutficient evidence to conclude that the new paint has a reflectometer reading higher than 20. Reject the null hypothesis.There is not sufticient evidence to condude that the new paint has a ref ectometer reading higher than 20. Do not reject the null hypothess.There is sultficent evidence to conclude that the new paint has a reflecto meter reading higher than 20 Reject the null hypothesis. There is sufficient evidence to conclude that the new pait has a reflectometer reading higher than 20
Answer #1

H0: \mu = 20 v/s Ha : \mu > 20

a) z = 3.3 , \alpha = 0.05

p value = p ( z > 3.3 ) = 1 - p ( z \leq 3.3 ) = 1 -  0.9995 = 0.0005

Here p value < \alpha (0.05 )

Reject null hypothesis . There is sufficient evidence to conclude that new paint has a reflectometer reading higher than 20.

b) z = 1.6 , \alpha = 0.01

p value = p ( z > 1.6 ) = 1 - p ( z \leq 1.6 ) = 1 - 0.9452 = 0.0548

Here p value > \alpha .

Do not reject null hypothesis. There is not sufficient evidence to conclude that the new paint has a reflectometer reading higher than 20.

c) z = -0.5, \alpha = 0.10

p value = p ( z > -0.5 ) = 1 - p ( z \leq -0.5 ) = 1 - 0.3085 = 0.6915

Here p value > \alpha

Do not reject null hypothesis. There is not sufficient evidence to conclude that the new paint has a reflectometer reading higher than 20.